x=7y-6
x+7y=3
find soultion of the equation

Answer by stanbon(75887)

x=7y-6
x+7y=3
------
Substitute for "x" and solve for "y"::
7y-6 + 7y = 3
---------------
14y = 9
------
y = 9/14
-----------------------
Solve for "x":
x = 7(9/14)-6
-----
x = (63/14)-(84/14)
-----
x = -21/14
-----
x = -3/2
-------------------
Cheers,
Stan H.